# Interaction-Augmented Instruction: Modeling the Synergy of Prompts and Interactions in Human-GenAI Collaboration

### v1

@@ 4.1. Revisit the Corpus
  • Ordered relations. Relations are numbered to encode the workflow order; relations that occur concurrently share the same index.
@@ 4.2. Atomic Paradigms
Using the IAI model as an analytic lens, we abstracted 12 recurring atomic paradigms (Table 2) from our corpus. We organize these paradigms along two orthogonal dimensions based on our model: (1) interaction timing: whether interaction (I) occurs before or after invoking GenAI (G); and (2) user resources: whether the user begins interacting with GenAI (G) when they have no artifact (A) at hand (prompt-only) or are in the artifact-grounded situation. The timing dimension roughly tracks intent clarity and control locus: pre-invocation interactions are typical when users can specify constraints up front, whereas post-invocation interactions support exploratory or ambiguous goals via mixed-initiative refinement. The resource dimension separates workflows where intent must be expressed solely in natural language from those where an existing artifact can be selected, annotated, or structured to ground the instruction. Below we analyze each of the four resulting classes as a coherent family of paradigms, and show how paradigm graph differences map to concrete tool behaviors via brief case comparisons.
@@ 4.2.1. Pre-invocation, Prompt-only: Structuring and Enhancing Prompts Before GenAI Invocation
This class groups paradigms (Table 2 P1-P3) where users begin with only prompts and seek to specify intent through pre-invocation refinement. Users typically have a clear goal and employ interaction to organize, extend, or transform textual instructions so that the ensuing generation aligns with their intent more precisely. For instance, a writer may elaborate a draft prompt with inline selection (OpenAI, 2024), or a programmer may decompose the prompt for a coding task into structured prompts for subtasks (Feng et al., 2024).
In all three paradigms (Fig. 3 P1-P3), humans write prompts (H $\rightarrow$ T) and GenAI produces artifacts (G $\rightarrow$ A), yet they diverge in how to enhance the original natural language prompts (i.e., refine T or construct Aug). The first paradigm, interactive prompt enhancement (P1), keeps the representation within the boundaries of natural language. User interactions just refine the wording of T (H $\rightarrow$ I $\rightarrow$ T), which is then executed directly. Dreamsheets (Almeda et al., 2024), for instance, enables users to rapidly compose text prompt variations with a set of keywords through spreadsheet-like interactions for exploratory image generation (Fig. 3-P1). While effective for quick iteration, it remains limited to text and cannot capture richer logical structures. In interactive prompt organization (P2), interactions introduce additional structure beyond natural language. The enriched representation, encoded as augmented instruction, integrates hierarchical or compositional logic before being passed to GenAI (H $\rightarrow$ I $\rightarrow$ T $\rightarrow$ Aug). CoLadder (Yen et al., 2024) illustrates this by arranging prompts into a tree of subtasks (Fig. 3-P2), while PromptChainer (Wu et al., 2022) links prompts in sequential chains. Yet even structured text has some limits; it presumes all user intent can be expressed linguistically. This gap is addressed by the third paradigm. Interaction as instruction (P3) encodes user actions themselves as operative intent, introducing non-linguistic signals in addition to text (H $\rightarrow$ I $\rightarrow$ Aug; T $\rightarrow$ Aug). SketchFlex (Lin et al., 2025), for example, interprets freehand sketches on a canvas as executable instructions combined with prompts to guide expressive image generation (Fig. 3-P3).
@@
Figure 3. Examples of pre-invocation paradigms, including prompt-only (P1-P3) and artifact-grounded (P4): (P1) Interactive prompt enhancement (Almeda et al., 2024); (P2) Interactive prompt organization (Yen et al., 2024); (P3) Interaction as instruction (Lin et al., 2025); (P4) Artifact as instruction (Masson et al., 2024).
@@
This class (Table 2 P4) captures paradigms where users start with an artifact and specify intent through pre-invocation manipulation (e.g., selection, brushing, sketching). It is also among the most common paradigms in existing systems. Typical scenarios include selecting a chart element to query (Masson et al., 2024), brushing an image region for editing (Liu et al., 2025), or highlighting code for debugging (Anysphere, 2023). Here, interactions ground the prompt in concrete referents before generation.
In the artifact as instruction paradigm (P4), humans interact with artifacts to encapsulate partial or entire its content into augmented instruction (H $\rightarrow$ I $\rightarrow$ A $\rightarrow$ Aug; T $\rightarrow$ Aug). Systems such as DirectGPT (Masson et al., 2024) and MagicQuill (Liu et al., 2025) demonstrate this approach by passing dragged or brushed elements as precise constraints (Fig. 3-P4). At the first glance at DirectGPT, it might be confused with SketchFlex (Fig. 3-P3) as they both create visual augmented instruction with interactions for GenAI to operate on artifacts. However, with our paradigm graphs, it is easy to notice that the key difference lies in how augmented instruction is composed: DirectGPT, as a representative of the artifact as instruction paradigm, includes raw artifact segments, while SketchFlex, which belongs to the interaction as instruction paradigm category, encodes non-linguistic signals with interactions (H $\rightarrow$ I $\rightarrow$ Aug $\rightarrow$ G $\rightarrow$ A). This comparison highlights the key difference between the two paradigms: whether the interaction carries intent solely without embedding artifact.
The design difference shapes how intent ambiguity or unclarity is resolved. The artifact as instruction paradigm (P4) mitigates referential ambiguity by passing raw content to GenAI, whereas the interaction as instruction paradigm (P3) reduces descriptive effort by letting interactions themselves encode intent. For editing and debugging tasks, pre-invocation artifact-grounded interactions are particularly effective as they tie model operations to specific referents.
Figure 4. Examples of post-invocation, prompt-only paradigms: (P5) AI-driven prompt suggestion (Angert et al., 2023); (P6) AI-driven prompt decomposition (Cai et al., 2024); (P7) Generative prompt control widgets (Wang et al., 2024a); (P8) Generative artifact control widgets (Vaithilingam et al., 2024).
@@ 4.2.3. Post-invocation, Prompt-only: Iterative Prompt Steering After GenAI Invocation
This class (Table 2 P5-P8) captures cases where users specify intent solely through prompts, but do so after an initial GenAI invocation. Unlike pre-invocation paradigms, where intent is clarified before execution, here the process is fundamentally iterative: the first prompt triggers model output, and subsequent refinement occurs in response to what GenAI returns. Because each atomic paradigm centers on a single GenAI role, these workflows often do not emphasize artifact creation, but rather focus on AI-assisted prompt steering and iterative negotiation of intent. This class arises when users’ initial instructions are vague, exploratory, or underspecified, and GenAI takes an active role in shaping subsequent prompts. Rather than users knowing exactly what they want, the system helps steer the process step by step.
@@
The first three paradigms in this class (Fig. 4 P5-P7) share the basic flow H $\rightarrow$ T $\rightarrow$ G (human writes text prompts to GenAI) and G $\rightarrow$ I (GenAI initiates interactions) but diverge in how AI-initiated interactions reshape subsequent instructions. In AI-driven prompt suggestion (P5), GenAI generates candidate refinements or alternative text prompts that users can adopt or edit (H $\rightarrow$ I $\rightarrow$ T), focusing on direct modification of natural language. Spellburst (Angert et al., 2023), for example, generates multiple auto-completed prompt suggestions by AI models. Then users can select one to guide subsequent calls (Fig. 4-P5). Similar to the extension from P1 to P2 (see Sec. 4.2.1), simple prompt refinement or suggestion by AI can hardly facilitate the needs for more structured intent communication. In AI-driven prompt decomposition (P6), GenAI externalizes its internal interpretation of a high-level prompt into a structured form such as a graph or task tree. Users then manipulate this representation directly (H $\rightarrow$ I $\rightarrow$ Aug), fragmenting, reordering, or parameterizing subtasks. Low-code LLM (Cai et al., 2024) and a recent work NeuroSync (Zhang et al., 2025), exemplify this approach by visualizing inferred code-generation plans as editable graphs (Fig. 4-P6). These paradigms still operate primarily through text manipulation, finer-grained specifications remain difficult to express.
@@
Furthermore, generative prompt control widgets (P7) extend prompts beyond natural language by having GenAI synthesize interactive controls. PromptCharm (Wang et al., 2024a), for instance, generates sliders tied to text spans, enabling users to adjust their relative weights within a text-to-image prompt (Fig. 4-P7). A related but more hybrid paradigm is generative artifact control widgets (P8), where GenAI produces both an artifact and associated widgets that persist for subsequent manipulation. DynaVis (Vaithilingam et al., 2024), for example, augments visualization-oriented natural language interfaces with dynamic controls that let users iteratively adjust and replay edits with instant feedback (Fig. 4-P8). While P7 and P8 both involve model-generated widgets, their scope differs: P7 parameterizes prompts, extending intent specification, whereas P8 binds controls to artifacts, turning outputs into malleable, persistent interaction surfaces. This distinction demonstrates the discriminative power of the IAI model and connects directly to recent work on malleable UIs and GenUI (Min et al., 2025), suggesting a trajectory where GenAI acts not only as a content generator but also as a co-designer of interfaces through which users iteratively shape intent.
@@
Figure 5. Examples of post-invocation, artifact-grounded paradigms: (P9) Artifact to organized instruction (Zhang et al., 2023); (P10) Artifact to multimodal instruction (Singh et al., 2024); (P11) Artifact-driven prompt enhancement (Chen et al., 2023); (P12) Interactive artifact refinement (Tang et al., 2024).
@@
This class (Table 2, P9–P12) captures workflows in which an artifact—either produced by a prior GenAI call or provided directly by the user—serves as the anchor for subsequent interaction. Rather than relying solely on textual prompts, GenAI inspects the artifact, initiates interactions, and elicits user feedback to clarify or refine intent. Typical tasks include iteratively editing an image, interrogating a visualization, or correcting model-produced code.
The common structure is artifact-driven (A $\rightarrow$ G), with GenAI initiating interactions (G $\rightarrow$ I) that engage human responses (H $\rightarrow$ I). The four paradigms differ in how these interactions shape the construction of follow-up instructions. In artifact to organized instruction (P9), GenAI analyzes the artifact, generates a set of candidate operations or reformulations, and presents choices that users select to compose a structured textual prompt. For example, VISAR (Zhang et al., 2023) proposes expansion options for a selected paragraph; the user’s choices produce a structured prompt tree for subsequent text generation (Fig. 5-P9). Building on this, artifact to multimodal instruction (P10) enriches the workflow by incorporating artifact itself. Here, GenAI extracts salient features from the artifact and asks users to tag or select them; the results are encoded into multimodal augmented instruction that combines artifact content with user intent as prompt. FigurA11y (Singh et al., 2024), for instance, extracts figure components and lets users link them to accessibility guidelines, producing a multimodal instruction (Fig. 5-P10). Moving beyond selection, artifact-driven prompt enhancement (P11) often positions the artifact as contextual grounding for semantic refinement. GenAI proposes contextual rules or mappings, which users confirm or modify in situ with interactions. LangAware (Chen et al., 2023) illustrates this by connecting low-level sensor signals to high-level contexts and enabling users to interactively combine contextual information with the original prompt before final execution (Fig. 5-P11). Finally, interactive artifact refinement (P12) emphasizes iterative analysis and correction. GenAI inspects the artifact in response to a prompt, surfaces candidate elements or annotations, and invites users to inspect, correct, or refine them. PDFChatAnnotator (Tang et al., 2024), for example, extracts information from PDFs and lets users guide annotation corrections interactively (Fig. 5-P12). Unlike prior paradigms, this focuses less on instruction construction and more on artifact-centered troubleshooting and refinement.

### v2

@@ 4.1. Revisit Corpora
  • Ordered relations. Relations are numbered to encode the workflow order; relations that occur concurrently share the same index number.
@@
To derive recurring paradigms, we canonicalized tool-specific labels into the standardized entities and relations and normalized away incidental UI differences that do not affect information flow. We then grouped the resulting atomic graphs by their topological structure to identify recurring interaction patterns. For each topological class, we generated a paradigm name, a canonical graph, a concise description, and a mapping to all systems exhibiting that structure. All tool metadata, annotation data, and grouping information are released to support replication222 The annotation results: https://interaction-augmented-instruction.github.io/.
Table 2. Taxonomy of atomic interaction paradigms in human-GenAI communication. The tools used to demonstrate corresponding paradigms in Sec. 4.2 are shown in bold.
Interaction Timing User
Resources
Paradigm Name Paradigm Description Atomic Paradigm Graph Example Tools
Interaction Before Calling GenAI Prompt-only (no artifact at hand) P1. Interactive Prompt Enhancement Human selects parts of drafted prompts for GenAI to refine or expand content. [Uncaptioned image] DreamSheets (Almeda et al., 2024), Promptify (Brade et al., 2023)
P2. Interactive Prompt Organization Human organize multiple text prompts into structured formats (e.g., tree) to fit the application. [Uncaptioned image] CoLadder (Yen et al., 2024), CoPrompt (Feng et al., 2024), PromptChainer (Wu et al., 2022), HoT (Nguyen et al., 2025), Intent Tagging (Gmeiner et al., 2025)
P3. Interaction as Part of Instruction Human’s interactions outside artifacts are included with prompts for GenAI to operate on artifacts. [Uncaptioned image] SketchFlex (Lin et al., 2025), InstructPipe (Zhou et al., 2025), Kim et al.  (Kim et al., 2023b)
Artifact-grounded P4. Referenced Artifact as Part of Instruction Human directly manipulates artifacts, which are combined with prompts as GenAI instructions. [Uncaptioned image] DirectGPT (Masson et al., 2024), Data Formulator (Wang et al., 2024a), ChatScratch (Chen et al., 2024), Qlarify (Fok et al., 2024), InternGPT (Liu et al., 2023b), MagicQuill (Liu et al., 2025), etc.
Interaction After
Calling GenAI
Prompt-only (no artifact at hand) P5. AI-driven Prompt Suggestion GenAI suggests new or extended prompts from the human’s initial input for human selection. [Uncaptioned image] Spellburst (Angert et al., 2023), Intent Tagging (Gmeiner et al., 2025), PromptCharm (Wang et al., 2024d), Bursztyn et al.  (Bursztyn et al., 2021), PromptCrafter (Baek et al., 2023), AiCommentator (Andrews et al., 2024), Sensecape (Suh et al., 2023), ExploreLLM (Ma et al., 2024), etc.
P6. AI-driven Prompt
Decomposition
GenAI restructures prompts into fine-grained, organized forms for interactive manipulation. [Uncaptioned image] Low-code LLM (Cai et al., 2024), NeuroSync (Zhang et al., 2025b), Liu et al.  (Liu et al., 2023a), SimStep (Kaputa et al., 2025), Fragments (Riche et al., 2025)
P7. Generative Prompt Control Widgets GenAI generates interactive widgets for fine-grained prompt control. [Uncaptioned image] PromptCharm (Wang et al., 2024d), DancingBoard (Chen et al., 2025b), Intent Tagging (Gmeiner et al., 2025)
P8. Generative Artifact Control Widgets GenAI generates widgets for humans to further manipulate or confirm artifact-related actions. [Uncaptioned image] DynaVis (Vaithilingam et al., 2024), Graphologue (Jiang et al., 2023), Luminate (Suh et al., 2024), Sparks (Gero et al., 2022), ConstitutionMaker (Petridis et al., 2024), ChatScratch (Chen et al., 2024), ScatterShot (Wu et al., 2023), etc.
Artifact-grounded P9. Artifact to Structured Instruction GenAI uses artifacts as starting points to generate structured prompts. [Uncaptioned image] VISAR (Zhang et al., 2023), Metaphorian (Kim et al., 2023a), Intent Tagging (Gmeiner et al., 2025)
P10. Artifact to Multimodal Instruction GenAI parses and integrates artifacts to construct instructions. [Uncaptioned image] FigurA11y (Singh et al., 2024)
P11. Artifact-driven Prompt Enhancement GenAI suggests actions based on contextual requests integrating artifacts and prompts. [Uncaptioned image] LangAware (Chen et al., 2023)
P12. Interactive Artifact Refinement GenAI analyzes artifacts and initiates interactions based on prompts. [Uncaptioned image] PDFChatAnnotator (Tang et al., 2024)
Figure 4.   Examples of pre-invocation paradigms (Refer to caption indicates the key interaction design in the interface, and red line highlights the key differences among the paradigm graphs), including prompt-only (P1-P3) and artifact-grounded (P4): (P1) Interactive prompt enhancement (e.g., Dreamsheets (Almeda et al., 2024)); (P2) Interactive prompt organization (e.g., CoLadder (Yen et al., 2024)); (P3) Interaction as part of instruction (e.g., SketchFlex (Lin et al., 2025)); (P4) Referenced artifact as part of instruction (e.g., DirectGPT (Masson et al., 2024)).
@@ 4.2. Atomic Paradigms
Using the IAI model as an analytic lens, we identified 12 recurring atomic paradigms (Tab. 2) from our corpus. We organize these paradigms along two orthogonal dimensions based on our model: (1) interaction timing: whether interaction (I) occurs before or after invoking GenAI (G); and (2) user resources: whether the user begins interacting with GenAI (G) when they have no artifact (A) at hand (prompt-only) or are in the artifact-grounded situation. The timing dimension roughly tracks intent clarity and control locus: pre-invocation interactions are typical when users can specify constraints up front, whereas post-invocation interactions support exploratory or ambiguous goals via mixed-initiative refinement. The user resource dimension separates workflows where intent must be expressed solely in natural language from those where an existing artifact can be selected, annotated, or structured to ground the natural language instruction.
These dimensions yield four paradigm classes with characteristic graph topologies and consistent design applicability. Atomic paradigm graphs are designed to capture human–GenAI interaction logic and information flow, rather than to fully depict end-to-end system execution, and therefore do not necessarily terminate at the same node. For example, in pre-invocation paradigms, GenAI typically acts as a generator: users initiate interactions to construct an augmented instruction, which is then executed by GenAI, making the generated artifact a natural endpoint of the graph. In post-invocation paradigms, by contrast, GenAI often plays a clarifying or interpretive role: it inspects existing artifacts or prior generated results and initiates interactions to elicit user input, usually resulting in a refined instruction rather than a new artifact. Consistent with our annotation principles that each atomic paradigm graph contains only a single GenAI role, passing this refined instruction to a generator GenAI is represented as a separate atomic graph, and complete application workflows can be expressed through chains of such graphs. Paradigms that involve no interaction are outside our scope, as our focus is on characterizing interaction structures and information flow rather than all possible execution paths.
To better showcase the model’s descriptive and discriminative power, accordingly, the following subsections adopt a unified structure: we first describe the shared characteristics and typical cases of each paradigm class; we then analyze how similarities and differences in their paradigm graph structures map to concrete operations of each tool through representative comparisons (Fig. 4, Fig. 5, Fig. 6); and we conclude each subsection by articulating the key intra-class distinctions and the design implications they entail.
@@ 4.2.1. Pre-invocation, Prompt-only: Structuring and Enhancing Prompts Before GenAI Invocation
This class groups paradigms (Tab. 2 P1-P3) where users begin with only prompts and seek to specify intent through pre-invocation refinement. Users typically have a clear goal and employ interaction to organize, extend, or transform textual instructions so that the ensuing generation aligns with their intent more precisely. For instance, a writer may elaborate a draft prompt with inline selection (OpenAI, 2024), or a programmer may decompose the prompt for a coding task into structured prompts for subtasks (Feng et al., 2024).
In all three paradigms (Fig. 4 P1-P3), humans write prompts (H $\rightarrow$ T) and GenAI produces artifacts (G $\rightarrow$ A), yet they diverge in how to enhance the original natural language prompts (i.e., refine T or construct Aug). The first paradigm, interactive prompt enhancement (P1), keeps the representation within the boundaries of natural language. User interactions just refine the wording of T (H $\rightarrow$ I $\rightarrow$ T), which is then executed directly. Dreamsheets (Almeda et al., 2024), for instance, enables users to rapidly compose text prompt variations with a set of keywords through spreadsheet-like interactions for exploratory image generation   (Fig. 4-P1 and Fig. 2-b). While effective for quick iteration, it remains limited to text and cannot capture richer logical structures. In interactive prompt organization (P2), interactions introduce additional structure beyond natural language. The enriched representation, encoded as augmented instruction, integrates hierarchical or compositional logic before being passed to GenAI (H $\rightarrow$ I $\rightarrow$ T $\rightarrow$ Aug). CoLadder (Yen et al., 2024) illustrates this by arranging prompts into a tree of subtasks (Fig. 4-P2), while PromptChainer (Wu et al., 2022) links prompts in sequential chains. Yet even structured text has some limits; it presumes all user intent can be expressed linguistically. This gap is addressed by the third paradigm. Interaction as part of instruction (P3) encodes user actions themselves as operative intent, introducing non-linguistic signals in addition to text (H $\rightarrow$ I $\rightarrow$ Aug; T $\rightarrow$ Aug). SketchFlex (Lin et al., 2025), for example, interprets freehand sketches on a canvas as executable instructions combined with prompts to guide expressive image generation   (Fig. 4-P3 and Fig. 2-c).
@@
Figure 5.   Examples of post-invocation, prompt-only paradigms (Refer to caption indicates the key interaction design in the interface, and red line highlights the key differences among the paradigm graphs): (P5) AI-driven prompt suggestion (e.g., Spellburst (Angert et al., 2023)); (P6) AI-driven prompt decomposition (e.g., Low-code LLM (Cai et al., 2024)); (P7) Generative prompt control widgets (e.g., PromptCharm (Wang et al., 2024d)); (P8) Generative artifact control widgets (e.g., DynaVis (Vaithilingam et al., 2024)).
@@
This class (Tab. 2 P4) captures paradigms where users start with an artifact and specify intent through pre-invocation manipulation (e.g., selection, brushing, sketching). It is also among the most common paradigms in existing systems. Typical scenarios include selecting a chart element to query (Masson et al., 2024), brushing an image region for editing (Liu et al., 2025), or highlighting code for debugging (Anysphere, 2023). Here, interactions ground the prompt in concrete referents before generation.
In the referenced artifact as part of instruction paradigm (P4), humans interact with artifacts to encapsulate either partial or entire content into augmented instruction (H $\rightarrow$ I $\rightarrow$ A $\rightarrow$ Aug; T $\rightarrow$ Aug). Systems such as DirectGPT (Masson et al., 2024) and MagicQuill (Liu et al., 2025) demonstrate this approach by passing dragged or brushed elements as precise constraints   (Fig. 4-P4 and Fig. 2-d). At first glance, DirectGPT could be mistaken for SketchFlex   (Fig. 4-P3 and Fig. 2-c) as they both create visual augmented instruction with interactions for GenAI to operate on artifacts. However, with our paradigm graphs, it is easy to notice that the key difference lies in how augmented instruction is composed: DirectGPT, as a representative of the referenced artifact as part of instruction paradigm, includes raw artifact segments, while SketchFlex, which belongs to the interaction as part of instruction paradigm category, encodes non-linguistic signals with interactions (H $\rightarrow$ I $\rightarrow$ Aug $\rightarrow$ G $\rightarrow$ A). This comparison highlights the key difference between the two paradigms: whether the interaction carries intent solely without embedding artifact.
The design difference shapes how intent ambiguity or unclarity is resolved. The referenced artifact as part of instruction paradigm (P4) mitigates referential ambiguity by passing raw content to GenAI, whereas the interaction as part of instruction paradigm (P3) reduces descriptive effort by letting interactions themselves encode intent. For editing and debugging tasks, pre-invocation artifact-grounded interactions are particularly effective as they tie model operations to specific referents.
@@ 4.2.3. Post-invocation, Prompt-only: Iterative Prompt Steering After GenAI Invocation
This class (Tab. 2 P5-P8) captures cases where users specify intent solely through prompts, but do so after an initial GenAI invocation. Unlike pre-invocation paradigms, where intent is clarified before execution, here the process is fundamentally iterative: the first prompt triggers model output, and subsequent refinement occurs in response to what GenAI returns. Because each atomic paradigm centers on a single GenAI role, these workflows often do not emphasize artifact creation, but rather focus on AI-assisted prompt steering and iterative negotiation of intent. This class arises when users’ initial instructions are vague, exploratory, or underspecified, and GenAI takes an active role in shaping subsequent prompts. Rather than users knowing exactly what they want, the system helps steer the process step by step.
@@
The first three paradigms in this class (Fig. 5 P5-P7) share the basic flow H $\rightarrow$ T $\rightarrow$ G (human writes text prompts to GenAI) and G $\rightarrow$ I (GenAI initiates interactions) but diverge in how AI-initiated interactions reshape subsequent instructions. In AI-driven prompt suggestion (P5), GenAI generates candidate refinements or alternative text prompts that users can adopt or edit (H $\rightarrow$ I $\rightarrow$ T), focusing on direct modification of natural language. Spellburst (Angert et al., 2023), for example, generates multiple auto-completed prompt suggestions by AI models. Then users can select one to guide subsequent calls (Fig. 5-P5). Similar to the extension from P1 to P2 (see Sec. 4.2.1), simple prompt refinement or suggestion by AI can hardly facilitate the needs for more structured intent communication. In AI-driven prompt decomposition (P6), GenAI externalizes its internal interpretation of a high-level prompt into a structured form such as a graph or task tree. Users then manipulate this representation directly (H $\rightarrow$ I $\rightarrow$ Aug), fragmenting, reordering, or parameterizing subtasks. Low-code LLM (Cai et al., 2024) and a recent work NeuroSync (Zhang et al., 2025b), exemplify this approach by visualizing inferred code-generation plans as editable graphs (Fig. 5-P6). These paradigms still operate primarily through text manipulation, but finer-grained specifications remain difficult to express.
@@
Furthermore, generative prompt control widgets (P7) extend prompts beyond natural language by having GenAI synthesize interactive controls. PromptCharm (Wang et al., 2024d), for instance, generates sliders tied to text spans, enabling users to adjust their relative weights within a text-to-image prompt   (Fig. 4-P7 and Fig. 2-e). A related but more hybrid paradigm is generative artifact control widgets (P8), where GenAI produces both an artifact and associated widgets that persist for subsequent manipulation. DynaVis (Vaithilingam et al., 2024), for example, augments visualization-oriented natural language interfaces with dynamic controls that let users iteratively adjust and replay edits with instant feedback (Fig. 5-P8). While P7 and P8 both involve model-generated widgets, their scope differs: P7 parameterizes prompts, extending intent specification, whereas P8 binds controls to artifacts, turning outputs into malleable, persistent interaction surfaces. This distinction demonstrates the discriminative power of the IAI model and connects directly to recent work on malleable UIs and GenUI (Min et al., 2025), suggesting a trajectory where GenAI acts not only as a content generator but also as a co-designer of interfaces through which users iteratively shape intent.
Figure 6.   Examples of post-invocation, artifact-grounded paradigms (Refer to caption indicates the key interaction design in the interface, and red line highlights the key differences among the paradigm graphs): (P9) Artifact to structured instruction (e.g., VISAR (Zhang et al., 2023)); (P10) Artifact to multimodal instruction (e.g., FigurA11y (Singh et al., 2024)); (P11) Artifact-driven prompt enhancement (e.g., LangAware (Chen et al., 2023)); (P12) Interactive artifact refinement (e.g., PDFChatAnnotator (Tang et al., 2024)).
@@
This class (Tab. 2, P9–P12) captures workflows in which an artifact—either produced by a prior GenAI call or provided directly by the user—serves as the anchor for subsequent interaction. Rather than relying solely on textual prompts, GenAI inspects the artifact, initiates interactions, and elicits user feedback to clarify or refine intent. Typical tasks include iteratively editing an image, interrogating a visualization, or correcting model-produced code.
The common structure is artifact-driven (A $\rightarrow$ G), with GenAI initiating interactions (G $\rightarrow$ I) that engage human responses (H $\rightarrow$ I). The four paradigms differ in how these interactions shape the construction of follow-up instructions. In artifact to structured instruction (P9), GenAI analyzes the artifact, generates a set of candidate operations or reformulations, and presents choices that users select to compose a structured textual prompt. For example, VISAR (Zhang et al., 2023) proposes expansion options for a selected paragraph; the user’s choices produce a structured prompt tree for subsequent text generation   (Fig. 4-P9 and Fig. 2-f). Building on this, artifact to multimodal instruction (P10) enriches the workflow by incorporating artifact itself. Here, GenAI extracts salient features from the artifact and asks users to tag or select them; the results are encoded into multimodal augmented instruction that combines artifact content with user intent as prompt. FigurA11y (Singh et al., 2024), for instance, extracts figure components and lets users link them to accessibility guidelines, producing a multimodal instruction (Fig. 6-P10). Moving beyond selection, artifact-driven prompt enhancement (P11) often positions the artifact as contextual grounding for semantic refinement. GenAI proposes contextual rules or mappings, which users confirm or modify in situ with interactions. LangAware (Chen et al., 2023) illustrates this by connecting low-level sensor signals to high-level contexts and enabling users to interactively combine contextual information with the original prompt before final execution (Fig. 6-P11). Finally, interactive artifact refinement (P12) emphasizes iterative analysis and correction. GenAI inspects the artifact in response to a prompt, surfaces candidate elements or annotations, and invites users to inspect, correct, or refine them. PDFChatAnnotator (Tang et al., 2024), for example, extracts information from PDFs and lets users guide annotation corrections interactively (Fig. 6-P12). Unlike prior paradigms, this focuses less on instruction construction and more on artifact-centered troubleshooting and refinement.
@@
Table 3.   Operationalizing Descriptive and Discriminative Power: Design cheat sheet mapping common design situations to recommended paradigms.
Design situation (intent formulation) IAI-model rationale Paradigm
Fast, lightweight iterations on text Interaction only edits text (H→I→T), keeping instruction purely textual. P1
Need structured intent or task decomposition Interaction adds explicit structure to text (H→I→T→Aug). P2
Need non-linguistic or spatial expressivity Interaction supplements instruction (H→I→Aug, T→Aug) beyond text. P3
Reduce referential ambiguity via artifacts Artifact content grounds intent (I→A→Aug, T→Aug) for precise reference. P4
GenAI-assisted exploration after an initial output GenAI proposes refinements (G→I); users adjust via light edits (H→I→T). P5
Need transparent intermediate reasoning GenAI exposes editable decomposition (G→I) for human adjustment (H→I→Aug). P6
Need parameter-level control without editing text GenAI initiates parameter widgets (G→I) that augment prompt (H→I→Aug, T→Aug). P7
Need persistent, artifact-bound controls for iteration Generated widgets (G→I) attach to artifacts, enabling repeated edits (H→I→A). P8
Have artifacts but need help forming follow-up instructions GenAI analyzes the artifact to propose structured follow-ups (A→G→I, H→I→T→Aug). P9
Need multimodal follow-up instructions grounded in artifacts GenAI derives multimodal instruction candidates from artifacts (A→Aug). P10
Need to integrate artifact-derived suggestions into prompts GenAI proposes artifact-based actions (A→G→I) that supply prompts (H→I→Aug, T→Aug). P11
Need explicit control over how operations apply to artifacts GenAI interprets artifacts to initiate grounded interactions (A→G→I) on artifacts (H→I→A). P12